If the supply and demand functions are given by  p=20e0.4Q and p=100e-0.2Q, respectively, find the equilibrium price and quantity, and calculate the consumer’s and producer’s surplus.


1
Expert's answer
2021-07-26T23:56:02-0400

Given,

Supply function:

"p=20e^{\n\n0.4Q}"

Supply function:

"p=100e^\n\n{\u22120.2Q}"

Note: the values will be in approx as the exponential function is given.

At equilibrium, quantity demanded is equal to the quantity supplied:

"20e^

{0.4Q}

=100e^

{−0.2Q}\\

\frac{e^

{0.4Q}}{

e^

{−0.2Q}}



=5

\\e^

{0.4Q+0.2Q}

=5\\

e^

{0.6Q}

=5\\

0.6Q=ln5 (e^

{x}

=y⇒x=ln(y))\\

0.6Q=1.6094\\

Q=\frac{1.6094}

{0.6}\\"

"Q=2.68 =3"  (nearest integer )

Substitute Q=2.68 in the supply function for price

"p=20\u00d7e^\n{\n\n0.4\u00d72.68}"

"p=20\u00d72.9212"

"p=58.424 = 58" (nearest integer)


Note: nearest integer has been taken for easy calculation.

Equilibrium price = 58 

Equilibrium quantity = 3

The formula for consumer surplus (CS) is given below:

"CS=∫^Q_0p(Q)dQ−p×Q (p(Q)=Demand \space function)\\CS=∫^{3}_0(100e^{−0.2Q})dQ−58×3\\CS=100[\frac {e^{−0.2Q}}{−0.2}]^3_0−174\\CS=\frac{100}{−0.2}[e^{−0.2×3}−e^{−0.2×0}]−174\\CS=−500[0.549−1]−174\\CS=−274.5+500−174\\CS=51.5 \space approx"

The value of consumer surplus is 51.5 approx.

The formula for producer surplus (PS) is given below:

"PS=p×Q−∫^Q_0p(Q)dQ (p(Q)=Supply \space function)\\PS=58×3−∫^3_0(20e^{0.4Q})dQ\\PS=174−20[\frac{e^{0.4Q}}{0.4}]^3_0\\PS=174−\frac{20}{0.4}[e^{0.4×3}−e^{0.4×0}]\\PS=174−50[3.320−1]\\PS=174−166+50\\PS=58

The value of pro"


The value of producer surplus is 58 approx.